OpenStudy (anonymous):

Determine whether the function is linear. Then find the evaluate for the given value of x. f(x)=x^2-3;f(2)

OpenStudy (anonymous):

The function is not linear, because it contains x^2. f(2)=2^2-3, so f(2) = 1

OpenStudy (anonymous):

For the answer at the end, do I do anything with the 2 that's next to the f or do I always have to leave it there?

OpenStudy (anonymous):

Hello?

OpenStudy (anonymous):

f(x) just means "the function of x", and in this case, the f(x) = x^2 -3. f(2) means "function of 2", so all you do is stick 2 into your function. You leave f(2) as it is.

OpenStudy (anonymous):

Ok, thank you!

OpenStudy (anonymous):

f(x)=x^2-3; this is a quadratic equation not a linear equation....f(2)=2^2 -3=4-3=1